OpenStudy (anonymous):

Please HELP!! Find all the solutions of the given equation: csc θ − cot θ = sin θ I've got the cosθ=0 and cosθ=1

OpenStudy (anonymous):

csc θ − cot θ = sin θ (1-cosθ )/sin θ = sin θ 1-cosθ= sin^2 θ 1-cosθ = 1- cos^2 θ (1-cosθ) = (1-cosθ)*(1+cosθ) (1-cosθ)*(1+cosθ-1)=0 so cosθ=0 and cosθ=1 which u have got is right

OpenStudy (anonymous):

Now cosθ=0 whenever θ is an odd integral multiple of pi/2 and cosθ=1 whenever θ is an even integral multiple of pi

OpenStudy (anonymous):

so the θ=pi/2+2kpi and θ=pi+2kpi? @matricked

OpenStudy (anonymous):

θ=2kpi +- pi/2 (or θ=pi/2+2kpi ) and θ=2kpi